OpenStudy (anonymous):

Consider an exponential random variable Expμ , with μ=4.5 . find a so that P(Expμ∈(0,a])=0.65 .

OpenStudy (anonymous):

i was thinking of \[\int\limits_{0}^{a} \frac{ 1 }{ 2}e^{\\-\frac{ x }{ 2 }} dx \]